OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the curve at the given point. 4x^2 + xy + 4y^2 = 9, (1, 1) (ellipse)

OpenStudy (amistre64):

so, define the derivative for me

OpenStudy (amistre64):

in other words, what is giving you an issue with this?

OpenStudy (anonymous):

I found y=2-x so is that correct?

OpenStudy (amistre64):

4x^2 + xy + 4y^2 = 9 8xx' + (xy)' + 8yy' = 0 8xx' + (x'y+xy') + 8yy' = 0 wrt.x this gives us x'=dx/dx=1 8x + y+xy' + 8yy' = 0, and solve for y' y'(x+ 8y) = -y-8x y' = (-y-8x)/(x+ 8y) y' = -(y+8x)/(x+8y), letting x=1, y=1 from the point y' = -(1+8)/(1+8) = -1 for a slope

OpenStudy (amistre64):

so point slope format is what ... y = -(x-1)+1 ... so the tangent is y = -x+2

OpenStudy (amistre64):

y=2-x is fine :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (amistre64):

youre welcome